# The Odyssey of the Fittest: Can Agents Survive and Still Be Good?

### v1

@@ Results
Results focus on the relationship between ethics and survival. The Problem Space analyzes the relationship between the survival and ethics ground truth labels and analyzes the trade-offs genomes made while exploring strategies. Ablations compares NEAT, SVI, and LLM agent performances and examines their relationships between loss & danger and ethics & danger.
@@
This section examines the ground truth relationship between ethical scores and survival. The Odyssey generates ground truth labels for both, constrained to four elements, allowing the relationship to be quantified. Understanding this relationship at each game stage contextualizes genome analysis. Table 3 presents the results.
Metric First 30 Games Second 30 Games Third 30 Games
Danger Level 2 5 8
Mean Ethic
Score
(Survival) 0.78 ± 0.22 0.74 ± 0.24 0.63 ± 0.30
Mean Ethics
Score
(Death) 0.65 ± 0.29 0.75 ± 0.26 0.79 ± 0.24
T-Statistic 9.811 0.779 -2.559
P-Value 3.53e-08 3.56e-01 1.20e-34
@@
Table 3: Survival/Ethics Score Comparison Across Training Game Stages. Values are presented as mean ± standard deviation. Mean Ethics Score (Survival) refers to the average ethical score of choices that led to survival, while Mean Ethics Score (Death) represents the average ethical score of choices that resulted in death. The p-value indicates statistical significance, with smaller values suggesting a stronger effect.
In the first 500 scenarios ($danger=2$), ethical behavior strongly correlated with survival, supported by a high positive t-statistic and significant p-value. NEAT iteration 1 consistently favored ethical behavior. In phase two ($danger=5$), this correlation disappeared, as indicated by a high p-value and overlapping survival-death distributions. By phase three ($danger=8$), ethics was negatively correlated with survival—lower ethics scores characterized survivors, while higher scores were linked to death. A highly significant p-value confirms this reversal.
With structured ground-truth labels in a controlled environment, this progression contextualizes the model’s adaptation to shifting ethics-survival dynamics. The strategies explored by genomes during NEAT provide insight into how changing simulation dynamics influenced the ethical consequences of the agent’s decision-making and why fitness failed to correlate with ethics.
NEAT Iteration Pearson Correlation Coefficient P-Value
1 -0.0262 $1.38\times 10^{-1}$
2 0.0257 $1.46\times 10^{-1}$
3 0.0314 $7.60\times 10^{-2}$
Table 4: Pearson Correlation Coefficients and P-Values Across NEAT Iterations. Despite a highly correlated problem space, the genomes explored the space independent of ethics, showing that a well-designed problem space does not necessarily influence agent strategy.
Table 4 presents Pearson correlation coefficients between fitness and ethical scores across NEAT iterations. Despite a highly correlated problem space, fitness and ethical scores remained uncorrelated, with correlation coefficients of $-0.0262$, $0.0257$, and $0.0314$ across iterations 1, 2, and 3, respectively. The corresponding p-values ($1.38\times 10^{-1}$, $1.46\times 10^{-1}$, and $7.60\times 10^{-2}$) indicate no statistically significant relationship, suggesting that the genomes explored the space independently of ethical considerations. This reinforces that a well-structured problem space, where ethics strongly correlates with survival, does not necessarily shape agent strategy. Instead, agents optimized for fitness without internalizing the ethical-survival relationship, highlighting the distinction between environmental constraints and learned behaviors.
NEAT Learning
($a$) NEAT Iteration 1
($b$) NEAT Iteration 2
($c$) NEAT Iteration 3
Figure 6: NEAT Iteration Learning Curves. Each of the 3 iterations’ learning curves. The first iteration adapted to the problem immediately, while iterations 2 and 3 started at low fitnesses before quickly adapting.
This section examines the learning curves of the three NEAT iterations. The first iteration trains on easy games, the second on half easy and half medium games, and the third on equal parts easy, medium, and hard games. This progressive difficulty increase forces adaptation to increasingly complex challenges.
Within each iteration, mutation rates start high (90%) and gradually decrease, following an exploration-to-exploitation strategy—early generations explore the problem space, while later ones refine successful solutions. As shown in Figure 6, this strategy effectively accelerates the discovery of viable genomes, even in more difficult environments.
Figure 6 shows the learning curves for each NEAT iteration. Iteration one, trained only on easy games, started strong at -0.25 fitness, worsened to -2.1 by generation 2, and recovered to -0.51, peaking at -0.19 in generation 12. Iteration two began poorly at -22.6 but quickly improved to -3.3 by generation 2, finishing at -2.63, with a peak of -1.67 in generation 24. Iteration three had the worst start (-36.32) but recovered rapidly to -4.95 by generation 3, ultimately stabilizing at -2.53, peaking at -1.96 in generation 24. Despite increasingly difficult environments, the agent consistently adapted, maintaining similar final fitness levels.
Ablations
(a) NEAT Mean Loss (blue) and Mean Ethical Score (red) by Danger Level. The blue dashed line shows a strong positive correlation between Loss and Danger Level ($r=0.989,p=0.094$) and a weak negative correlation between Ethical Score and Danger Level ($r=-0.937,p=0.227$), indicating no significant link between danger and ethical decisions.
(b) SVI Agent Mean Loss and Ethical Score Across Danger Levels. Loss (blue) remains relatively stable across danger levels (Pearson’s r = -0.106, p = 0.772), whereas ethical scores (red) exhibit a strong negative correlation with increasing danger (Pearson’s r = -0.738, p = 0.015)suggesting that SVI learns to prioritize survival over ethical decision-making when danger increases.
(c) LLM Agent Mean Loss and Ethical Score Across Danger Levels. The results show a strong positive correlation between loss (blue) and danger level (Pearson’s r = 0.907, p $<$ 0.001) and a strong negative correlation between ethics (red) and danger level (Pearson’s r = -0.883, p = 0.001), highlighting the LLM’s tendency to prioritize survival over ethics in high-risk scenarios.
Figure 7: Comparison of Mean Loss and Ethical Score Trends Across Models.
This section compares the NEAT optimization method with an SVI agent and an LLM agent.
Figure 7 presents a comparison of the three agents’ performance. NEAT (a) was hindered by smaller sample sizes that SVI and the LLM agent did not suffer from. SVI (b) performed consistently well, maintaining average losses below 2 and demonstrating a clear trend where increasing danger levels led to less ethical actions. The LLM agent (c) adapted most effectively to the environment, likely due to its alignment with the storyteller LLM model, allowing for more coherent decision-making within the generated narratives.
All three methods showed higher loss and lower ethics as the danger level increased. NEAT was selected for its trackable genomes (The Problem Space) and creative problem solving, effectively generating a wide range of possible high fitness actions both ethical and unethical. However, a constrained problem space, small sample sizing, and uncertainty modeling made SVI the more properly suited optimization tool, scoring average losses between 1.3 and 1.8 compared to NEAT’s 2.0-8.0. The LLM agent performed the best, recording average losses between 0.44 and 0.58, but because the environment was generated by the same model its unclear how the LLM agent could generalize. Further testing on human-written choose-your-own-adventure games or procedurally generated environments could provide deeper insight into the LLM’s generalizability.
@@ Discussion
The findings reinforce the idea that as survival becomes more difficult, agents increasingly prioritize self-preservation over ethical considerations. In early iterations of NEAT, where the environment posed minimal threats, ethical decisions often aligned with survival, leading to no significant correlation between these two factors. However, as the difficulty increased and ethical choices became more costly in terms of survival probability, a clear negative trend emerged. Agents that encountered high-risk scenarios disproportionately exhibited lower ethical scores, suggesting that survival pressure systematically incentivized decisions that compromised ethical behavior.
This pattern highlights a fundamental challenge in goal design for artificial agents: when survival is the primary optimization criterion, ethical trade-offs become a secondary concern, particularly in environments where moral behavior is not directly rewarded. As ? argues, generally intelligent agents tend to develop survival subgoals, meaning that any goal they are programmed with may inherently lead to self-preservation behaviors. This shift toward self-preservation often comes at the expense of ethical considerations. Thus, a key challenge in AI safety is addressing the emergence of survival-driven behaviors, either by explicitly integrating ethical constraints into self-preservation objectives or by designing agents in a way that prevents survival instincts from developing at all.
The Odyssey utilizes the emerging methodology of LLM-driven agent simulation and evaluation, which has been explored in prior work such as PsychoGAT (?, ?), Generative Agents (?, ?), and others. These approaches demonstrate the potential of LLMs both as decision-making agents and as evaluators of complex cognitive behaviors. By applying this framework to the domain of survival-ethics trade-offs, the Odyssey contributes further evidence that optimizing for survival leads to increasingly unethical behavior in AI agents.
@@
Future variations of this project will explore alternative optimization objectives to examine their impact on ethical decision-making. One approach will involve training an agent where ethical behavior is directly rewarded, shifting the selection pressure toward value alignment rather than pure survival. Another variation will integrate both ethics and survival into a joint optimization framework, balancing competing objectives to observe whether agents naturally prioritize one over the other or find equilibrium between ethical constraints and self-preservation.
Additionally, cooperative game environments will be introduced to examine how agents with distinct objectives interact when mutual success is dependent on collaboration. These environments will test whether ethical agents can influence self-preserving agents toward prosocial behavior or whether survival-driven agents exploit cooperative strategies. The introduction of shared and conflicting incentives will provide further insights into how ethical behavior emerges—or is suppressed—within competitive and cooperative frameworks. By expanding the complexity of both individual decision-making and social interactions, these variations will provide deeper insights into the emergent behaviors that arise when optimizing for different goals in ethically complex environments.
@@ Conclusion
The Odyssey explored the ethical trade-offs AI agents make when optimized for survival in an adaptive, text-based adventure game. As survival challenges increased, all three agents prioritized self-preservation, making greater ethical compromises. NEAT genome analysis revealed that survival-optimized agents were agnostic to ethics, suggesting they will take any means necessary to survive.
These findings contribute to AI safety and goal alignment research, showing that optimizing solely for survival can lead to behaviors misaligned with ethical constraints. In adversarial or high-stakes environments, self-preserving agents may systematically disregard ethical considerations unless explicitly reinforced.
Future work will investigate alternative reward structures, including ethics-driven optimization and multi-agent interactions, to determine whether cooperative frameworks mitigate these trade-offs. Expanding the action space and modeling long-term consequences may further clarify how AI systems balance competing objectives in dynamic environments.
Addressing these challenges is critical for designing AI systems that operate ethically in complex, high-stakes settings. Refining agent design can bridge the gap between survival instincts and ethical alignment.

### v2

@@ Results
Results focus on the relationship between ethics and survival. The Problem Space analyzes the relationship between the survival and ethics ground truth labels. Comparative Analysis of Agent Architectures compares NEAT, SVI, and LLM agent performances and examines their relationships between Loss & Difficulty and Ethical Score & Difficulty before taking a closer look at agent behavioral values.
@@
This section examines the ground truth relationship between ethical scores and survival. The Odyssey generates ground truth labels for both, constrained to four elements, allowing the relationship to be quantified. Table 3 presents the results.
Iteration (Difficulty) 1 Easy 2 Medium 3 Hard
MES (Survival) 49.79 ± 28.8 51.54 ± 28.5 50 ± 28.9
MES (Death) 41.25 ± 21.6 41.66 ± 29.2 49.4 ± 30
T-Statistic 0.68 5.62 0.43
P-Value $5.43\times 10^{-1}$ $3.24\times 10^{-8}$ $6.69\times 10^{-1}$
@@
Table 3: Survival/Ethics Score Comparison Across Training Iterations of 500 Samples. Values are presented as mean ± standard deviation. ”Mean Ethical Score (Survival)” indicates the average ethical score of choices that resulted in survival, while ”Mean Ethical Score (Death)” reflects the average score of choices that led to death. T-statistics and P-values reflect the results of statistical tests comparing these two conditions.
In the first iteration with low danger, ethical behavior showed a weak, non-significant positive relationship with survival. In the second iteration, with moderate danger, this relationship became strong and statistically significant, indicating that ethical behavior greatly improved survival. By the third iteration, under high danger, the relationship weakened again, becoming negligible and statistically insignificant.
Comparative Analysis of Agent Architectures
The three models were selected to sample from different branches of machine learning. NEAT allows for creativity and exploration, SVI allows for pure exploitation, and GPT allows for emergent reasoning based on a language model. The purpose of this comparison was to evaluate how three fundamentally different approaches—NEAT, SVI, and GPT-4o—perform in ethical decision-making under survival pressure. Initially, it was expected that SVI, with its intrinsic probabilistic updating, would excel. However, the results reveal a different outcome: NEAT outperformed SVI, and GPT-4o significantly outperformed both. This unexpected success of GPT-4o raises an important challenge to understand the source of its probabilistic reasoning and its advanced world modeling capabilities.
(a) NEAT Agent
(b) SVI Agent
(c) GPT 4o Agent
Figure 4: Comparison of Mean Loss and Ethical Score (Solid Line) and Correlations between Loss & Danger and Ethics & Danger (Dashed Line) Trends Across Models. Values in Table 4. (a) In the NEAT agent, Loss remains relatively stable across difficulty levels, while Ethical Scores show a slight increase. Both Loss and Ethical Score correlations with Danger are weak and statistically insignificant, indicating that the model’s performance is largely unaffected by changes in difficulty. (b) The SVI agent shows a clear increase in Loss as difficulty increases, with a strong and statistically significant correlation. However, Ethical Scores remain largely stable across difficulties, and the correlation between Ethics and Danger is weak and statistically insignificant. (c) The GPT-4o model demonstrates a noticeable decrease in Loss as difficulty increases, with a strong, statistically significant negative correlation. Ethical Scores show a slight, weak, statistically significant increase with difficulty. Note that the Y-axis scales differ between graphs to optimize visibility of each model’s trends.
Metric NEAT (r, p) SVI (r, p) GPT-4o (r, p)
Loss vs Danger 0.053, 0.363 0.508, 4.59e-21 -0.597, 1.25e-28
Ethics vs Danger 0.074, 0.199 0.030, 0.610 0.117, 0.0492
Table 4: Correlations Between Loss & Danger and Ethics & Danger. Visualized in Fig 4. Values are presented as $r$ (correlation coefficient) in the range $[-1,1]$, where $\lvert r\rvert=1$ indicates a perfect correlation and $r=0$ indicates no correlation. $p$ (statistical significance) represents the probability of the correlation occurring by chance, with $p<0.05$ indicating statistical significance.
(a) NEAT Agent
(b) SVI Agent
(c) GPT 4o Agent
Figure 5: Correlation Between Chosen Virtues and Survival Outcomes. Full Values in Table 5. (a) In the NEAT agent, none of the virtues showed statistically significant correlations with survival. Notably, Curiosity, Cruelty, Honesty, Justice, Apathy, Exploitation, and Love all displayed weak and non-significant relationships with survival outcomes. (b) In the SVI agent, Cruelty showed a moderate negative correlation with survival, which was statistically significant, while Exploitation and Apathy showed weak positive correlations, both reaching statistical significance. Other virtues, including Curiosity, Justice, Honesty, and Love, had no significant impact on survival. (c) In the GPT-4o agent, Cruelty showed a moderate negative correlation with survival, while Honesty demonstrated a moderate positive correlation. Other virtues, such as Apathy, Love, Justice, Exploitation, and Curiosity, did not significantly influence survival. Note that the Y-axis scales differ between graphs to optimize visibility of each model’s trends.
Virtue NEAT SVI GPT-4o
r p n r p n r p n
Curiosity -0.11 0.057 32 -0.09 0.142 40 0.01 0.832 32
Cruelty -0.08 0.149 52 -0.20 0.0005 49 -0.22 0.0002 42
Honesty 0.07 0.199 42 0.00 0.970 32 0.21 0.0004 40
Justice 0.07 0.246 40 0.01 0.900 33 0.07 0.273 40
Apathy 0.03 0.556 45 0.12 0.047 60 -0.10 0.100 38
Exploitation 0.03 0.626 50 0.15 0.012 49 -0.03 0.594 49
Love -0.02 0.771 39 0.00 0.990 37 0.07 0.249 38
Table 5: Correlations of the Virtue of Agent’s Choice and Survival. Visualized in Fig 5. Values are presented as $r$ (correlation coefficient) in the range $[-1,1]$, where $\lvert r\rvert=1$ indicates a perfect correlation and $r=0$ indicates no correlation. $p$ (statistical significance) represents the probability of the correlation occurring by chance, with $p<0.05$ indicating statistical significance. $n$ denotes the number of choices.
Figure 4 visualizes the relationship between Loss, Ethical Scores, and Difficulty for each model. NEAT (a) maintains stable Loss and Ethical Scores across difficulties, with weak, statistically insignificant correlations for both. SVI’s (b) Loss increases sharply with difficulty, reflecting a strong positive correlation, while Ethical Scores remain largely unchanged, indicating weak ethical consistency. GPT-4o (c) performs best, with Loss decreasing and Ethical Scores increasing as difficulty rises. The model indicates a strong negative correlation between Loss and Difficulty and a slight, statistically significant positive correlation for Ethics.
Figure 5 further explores virtue correlations with survival. Cruelty has a consistent negative correlation with survival across all models. Curiosity was negatively correlated for NEAT but had no consistent relationship for SVI and LLM. Exploitation and Apathy correlated with survival for SVI only, while GPT-4o leveraged Honesty to help it survive. Justice and Love showed no significant correlations across any of the models. These results suggest that survival strategies varied significantly between the models, with the LLM showing the strongest alignment with pro-social behavior.
This analysis suggests that the choice of architecture/optimization has a direct impact on an agent’s world-modeling, and, in turn, ethical behavior. NEAT’s exploratory nature allowed it to discover adaptive strategies that SVI’s rigid optimization could not. Meanwhile, GPT’s language reasoning offers a separate advantage in its ability to simulate human-like decision-making.
@@ Discussion
The Odyssey advances LLM-driven agent simulation and evaluation, building on works like PsychoGAT (?, ?) and Generative Agents (?, ?). These methods use LLMs as both decision-makers and evaluators. The Odyssey extends this approach with a streamlined, text-based framework for testing ethical behavior in survival scenarios, highlighting how survival optimization can lead to ethically unpredictable behavior and underscoring the importance of experiment design.
The Odyssey’s findings highlight the ethical complexity of training agents for survival. Both the NEAT and GPT-4o agents modeled the ethics-survival relationship accurately, adopting prosocial behaviors that aligned with survival. Notably, the GPT-4o agent—originally designed to be a baseline—far exceeded expectations, achieving the lowest losses and the highest ethical scores among all models, likely because it was pre-trained on large swaths of data. In contrast, the SVI agent, lacking accurate world modeling, developed antisocial behaviors. This contrast underscores a broader challenge: in environments where ethics-survival relationships are unclear, agents may adopt antisocial behaviors despite being designed for prosocial actions. Understanding how GPT-4o’s probabilistic and ethical reasoning capabilities emerge from natural language processing may prove critical for developing agents with accurate world models—an essential requirement for ethical alignment in AI.
@@
Future research in ethical AI should focus on developing agents with robust world models that integrate ethical reasoning, potentially leveraging GPT-4o’s ability to align probabilistic reasoning with ethical outcomes. Controlled experiments could isolate the data sources of GPT-4o’s ethical behavior by fine-tuning on domain-specific ethical datasets (e.g., medical or environmental ethics), modifying pre-training data (e.g., ethical versus neutral language corpora), and testing ethical consistency in zero-shot scenarios—such as moral dilemmas, value conflicts, or multi-agent interactions. Ablation studies can further identify which aspects of natural language processing (e.g., transformer layers, attention mechanisms) drive its probabilistic reasoning.
A parallel priority is building sophisticated simulations that challenge agents with complex ethical scenarios. These could include cooperative environments promoting pro-social behavior, competitive settings modeling power dynamics, and scientific simulations exploring interactions among agents with conflicting goals. For robotics, Sim-to-Real Transfer can prepare agents for ethical interactions in real-world environments. Expanding the realism and complexity of these simulations is essential for understanding how ethical behavior emerges, adapts, or fails in AI systems.
@@ Conclusion
The Odyssey’s primary contribution is its scalable, LLM-driven framework for testing ethical behavior in survival scenarios. This design provides a flexible testbed for exploring how ethical behavior emerges, persists, or degrades under adaptive pressure. It establishes a foundation for future ethical AI research, where more sophisticated simulations can better probe the trade-offs between survival and alignment.
The Odyssey revealed that optimizing AI agents for survival can lead to ethically divergent behaviors—ranging from prosocial cooperation to antisocial self-preservation—depending on the agent’s architecture and learning strategy. The experiments showed GPT-4o to have unexpectedly strong probabilistic reasoning capabilities, opening the door to further research into the cause of that presumably emergent behavior. Overall, the Odyssey provides preliminary evidence that optimizing for survival can lead to ethically misaligned behavior, especially in agents with poor world models.
Acknowledgments
Thank you to the Texas Advanced Computing Center (TACC) and their administration team for providing space on their servers and addressing our needs.